Consider two risky investments with the following return distributions:

Probability

Return R1 (p.a.)

0.25

0.30

0.25

0.20

+12%

+4%

−5%

−8%

Expected return

μ1 = 1.3500%

Volatility

σ1 = 7.6176%

Probability

Return R2 (p.a.)

0.30

0.30

0.20

0.20

+10%

+8%

+3%

−15%

Expected return

Volatility

The correlation between the returns of the two investments is ρ12 = 0.9.

  1. (a) Calculate μ2 and σ2. (Read the Instructions carefully.)

  2. (b) Express the portfolio squared volatility σP2 in terms of w1, the weight of the first investment. (Round the coefficients to 6 decimals.)

  3. (c) Using your answer in part (b), find the weight w1 of the minimum-risk portfolio. Hence, find the expected return μP and volatility σP of the minimum-risk portfolio.
Answer #1
Expected Return =Mean Return =SUMof ((Probability)*(Return))
Variance of Return =Sum of(Probability* (Deviation ^2))
Deviation =Return -Mean Return
Standard Deviation of Return =Square Root of Variance of Return
ANALYSIS OF ASSET 1
p R1 A1=R1*P D1=R1-1.35 E1=(D1^2) F1=p*E1
Probability Return(%) Probability*Return(%) Deviation(%) Deviation Squared(%%) Probability*Deviation Squared(%%)
0.25 12 3 10.65 113.4225 28.355625
0.30 4 1.2 2.65 7.0225 2.10675
0.25 -5 -1.25 -6.35 40.3225 10.080625
0.20 -8 -1.6 -9.35 87.4225 17.4845
SUM 1.35 SUM 58.0275
Expected Return =Mean return 1.35 %
Variance of Return 58.0275 %%
Standard Deviation of Return =SQRT(58.0275)= 7.6176 %
ANALYSIS OF ASSET 2
p R2 A2=R2*p D2=R2-2.65 E2=(D2^2) F2=p*E2
Probability Return(%) Probability*Return(%) Deviation(%) Deviation Squared(%%) Probability*Deviation Squared(%%)
0.25 10 2.5 7.35 54.0225 13.505625
0.30 8 2.4 5.35 28.6225 8.58675
0.25 3 0.75 0.35 0.1225 0.030625
0.20 -15 -3 -17.65 311.5225 62.3045
SUM 2.65 SUM 84.4275
Expected Return =Mean return 2.65 %
Variance of Return 84.4275 %%
Standard Deviation of Return =SQRT(84.4275)= 9.1884 %
Covariance between Return of Asset 1 and asset 2=Correlation*Standard Deviation of asset 1*Standard Deviation of asset 2
Covariance between Return of Asset 1 and asset 2 62.9943 %% (0.9*7.6176*9.1884)